A new cream that advertises that it can reduce wrinkles and improve skin was subject to a recent study. A sample of 55 women over the age of 50 used the new cream for 6 months. Of those 55 women, 37 of them reported skin improvement(as judged by a dermatologist). Is this evidence that the cream will improve the skin of more than 40% of women over the age of 50? Test using α=0.01. (a) Test statistic: z= (b) Critical Value: z∗= (c) The final conclusion is A. We can reject the null hypothesis that p=0.4 and accept that p>0.4. That is, the cream can improve the skin of more than 40% of women over 50. B. There is not sufficient evidence to reject the null hypothesis that p=0.4. That is, there is not sufficient evidence to reject that the cream can improve the skin of more than 40% of women over 50.
Answer #1

a)

H0: p = 0.40

Ha: p > 0.40

Sample proportion \hat{p} = 37/55 = 0.6727

Test statistics

z = \hat{p} - p / sqrt( p ( 1 - p) / n)

= 0.6727 - 0.40 / sqrt( 0.40 * 0.60 / 55)

= 4.13

b)

Critical value Z* = 2.326

c)

Since test statistics z > 2.326, we have sufficient evidence to reject H0.

Conclusion -

We can reject the null hypothesis that p = 0.4 and accept that p > 0.4. That is, the cream can improve the

skin of more than 40% of women over 50.